OpenStudy (anonymous):

is X the solution of the equation ? x=-3 -3x^2=-27

OpenStudy (jdoe0001):

well... dunno so let's use the given "x" the given "x" is x = -3 so \(\bf x=-3\qquad \qquad -3x^2=-27\\ \textit{let us use }x = -3\\ \quad \\ -3x^2=-27\implies -3(-3)^2=-27\implies -3\cdot 9= -27\) what do you think? does the left-hand-side EQUALS the right-hand-side?